OpenStudy (anonymous):

If any tangent to the ellipse \frac{X^2}{a^2} + \frac{Y^2}{b^2}=1 \) intercepts equal length \(L\) on the axes then find \( L\) Sadly it took me more than 5 mints to solve this, how much do you take? ... Am I having intellectual atrophy?

OpenStudy (binary3i):

you just have to write the equation if the form Y=F(x) then find its derivative, equate the derivative equation to -1 and find your X which is your L

OpenStudy (amistre64):

what does "intecepts equal length" mean?

OpenStudy (binary3i):

intercepts of the tangent are equal, i guess.

OpenStudy (anonymous):

It is length of the intercept

OpenStudy (amistre64):

so essentially, (0,L) and (L,0) ??

OpenStudy (anonymous):

Yes exactly

OpenStudy (amistre64):

\[ \frac{X^2}{a^2} + \frac{Y^2}{b^2}=1 \] \[ \frac{2X}{a^2} + \frac{2Y}{b^2}Y'=0 \] \[ \frac{X}{a^2} + \frac{Y}{b^2}Y'=0 \] \[ \frac{Y}{b^2}Y'=-\frac{X}{a^2} \] \[ Y'=-\frac{b^2}{a^2}\frac{X}{Y} \] a slope of \(\pm\)45 degrees should create the desired results right?

OpenStudy (amistre64):

or when Y' = 1 or -1

OpenStudy (amistre64):

(x,y) = (a^2,b^2) or some signed version of it

OpenStudy (anonymous):

The answer is sqrt(a^2+b^2) but I dont know how to do it

OpenStudy (amistre64):

ah, so not just one point, but a locus of points ... is what that looks like to me

OpenStudy (amistre64):

create the equation of the tangent line\[y = -\frac{Xb^2}{Ya^2}(x-X)+Y\] let (x,y) -> (a,b) be any point on the ellipse \[b = -\frac{Xb^2}{Ya^2}(a-X)+Y\] and solve for Y

OpenStudy (amistre64):

if memory serves

OpenStudy (anonymous):

I did it but in vain

OpenStudy (amistre64):

\[b = -\frac{Xb^2}{Ya^2}(a-X)+Y\] \[b +\frac{Xb^2}{Ya^2}(a-X)=Y\] \[Ya^2b +Xb^2a-X^2b^2=Y^2a^2\] \[Xb^2a-X^2b^2=Y^2a^2-Ya^2b\] i should prolly check to see if ive remembered the gerenal tangent line setup correctly :)

OpenStudy (amistre64):

but as is, its starting to resemble a circle equation that you would need to complete the squares on

OpenStudy (amistre64):

the general setup for a tangent line from a derivative is: y = f'(a)(x-a) + f(a); for all points (a,f(a)) on f(x) our intercepts can then be defined as: when x=0, y = f'(a)a + f(a) and when y = 0, x = a - f(a)/f'(a) and you want them to be equal ....

OpenStudy (amistre64):

dropped a negative on that y intercept :/ -f'(a)a + f(a)

OpenStudy (amistre64):

-f'(a)a + f(a) = a - f(a)/f'(a) -f'^2(a)a + f'(a)f(a) = f'(a)a - f(a) a f'^2(a) - (f(a)-a) f'(a) - f(a) = 0 interesting, that turns out to be rather quadratic in general

OpenStudy (anonymous):

I got the answer